OpenStudy (nuccioreggie):

Write an expression that is equivalent to two over three (4x + 9). eight over threex + 6 eight over threex + 9 two over threex + 6 two over threex + 9

OpenStudy (anonymous):

\(\dfrac{2}{3} \times \dfrac{4}{1}\) Multiply 2 * 4 and 3 * 1

OpenStudy (nuccioreggie):

Its B right